[(sqrt(3+sqrt5)) - (sqrt(3-sqrt5))]^2 = 2
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

[(sqrt(3+sqrt5)) - (sqrt(3-sqrt5))]^2 = 2
-----
Take the sqrt of both sides to get:
sqrt(3+sqrt(5)) - sqrt(3-sqrt(5)) = sqrt(2)
-----
2.2882 - 0.874 = 1.414
1.414 = 1.414
---------
Cheers,
Stan H.
------------------

Answer by solver91311(24713)





Expand the binomial in the LHS



Collect like terms and calculate the product of conjugates in the center term of the LHS





John

My calculator said it, I believe it, that settles it